sara drived off going west at a speed of 40 miles per hour.Now 2 hrs later, Jay decides to go after her and takes off at a speed of 90miles per hour. How long after Sara left will it be when Jay catches her?

Answer by stanbon(75887)

Sara drove off going west at a speed of 40 miles per hour.
Now 2 hrs later, Jay decides to go after her and takes off at a speed of 90miles per hour.
How long after Sara left will it be when Jay catches her?
----------
Sara DATA:
rate = 40 mph ; time = t hrs; distance = rt = 40t miles
-------------
Jay DATA:
rate = 90 mph ; time = t-2; distance = rt = 90(t-2)
-------------
Equation:
distance = distance
40t = 90(t-2)
40t = 90t - 180
50t = 180
t = 18/5 = 3.6 hrs.
======================
Cheers,
Stan H.